OpenStudy (anonymous):

Evaluate -9(x – (-1)) ÷ 6 – x2 for x = -1. Please show all of your work.

OpenStudy (anonymous):

-9(x – (-1)) ÷ 6 – x^(2) ? = -9(x – (-1))/(6 – x^(2)) distribute the -9, remember -(-) = +

OpenStudy (anonymous):

It's 0. -1 for x makes everything in the numerator equal zero. Regardless of the denominator, your answer is zero.

OpenStudy (anonymous):

ok im confused

OpenStudy (anonymous):

In the numerator, you get 9*(-1 - (-1)). -1 - (-1) is zero. 9 times zero is zero. Zero divided by anything is zero.

OpenStudy (anonymous):

-9(x – (-1)) = -9(x + 1)) sub in x= -1 -9(-1 + 1) = -9(0) = 0 whenever you have a numerator of 0 the answer is 0 unless it is 0/0 then it is undefined, you cannot divide by zero

OpenStudy (anonymous):

cuddlepony explained it better

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

i got it now

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

0 = (1 + x)/2 notice when I try to find the roots for this 0*2 = ((1+x)/2)*2 0 = 1 + x Therefore it doesn't matter what is on the denominator, the root is always dependent on the numerator